OpenStudy (anonymous):

Did i find the inverse functio correctly? For y=-7log_6 (22x) My answer: 6^(x/7) / 2 iinverse of above

OpenStudy (anonymous):

I mean 2x not 22x

zepdrix (zepdrix):

Hmm I'm trying to figure out where your negative sign went.

zepdrix (zepdrix):

\[\Large\rm x=-7\log_6(2y)\]\[\Large\rm -\frac{x}{7}=\log_6(2y)\]You should end up with a negative in your exponent, ya? :)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Did i get it right, i think i forgot a negative

OpenStudy (anonymous):

I convert to exponential form right?

zepdrix (zepdrix):

Yah that seems like a good step right there.

zepdrix (zepdrix):

The rest of your solution looks correct though. \[\Large\rm y^{-1}(x)=\frac{6^{(-x/7)}}{2}\] Just missing the negative :)

OpenStudy (anonymous):

Yay thanks. Can you also check: y=log_3 (2^y-10) Answer is 2^(3x+10) Sorry. I dont have answer key, and im not sure if im doing these corrctly xD

zepdrix (zepdrix):

k sec :)

zepdrix (zepdrix):

\[\Large\rm y=\log_3(2^y-10)\]Is this what the problem looks like? The -10 is not in the 2 exponent right?

zepdrix (zepdrix):

Woops that should be 2^x hehe

zepdrix (zepdrix):

oh you made the same typo, that's funny lol

zepdrix (zepdrix):

\[\Large\rm x=\log_3(2^y-10)\]Converting to exponential,\[\Large\rm 3^x=2^y-10\]Add 10,\[\Large\rm 3^x+10=2^y\]Then I think you made a boo boo right here. We need to undo this exponential base 2. We do that by log_2'ing each side.

zepdrix (zepdrix):

\[\Large\rm \log_2(3^x+10)=\log_2(2^y)\]

OpenStudy (anonymous):

Oh ok ty, you solve and set equal